a solid ball of radius R and mass m is pushed with an initial velocity of 3 m / s onto an inclined plane of 30 degrees with a kinetic friction coefficient of 0.2 and a coefficient of static friction of 0.3. When climbing the inclined plane, the ball always rolls and never slips. The maximum height the ball can reach is approx
1
Expert's answer
2020-08-17T08:58:03-0400

According to the law of conservation of energy, the change in kinetic and potential energy is equal to the work done by friction force.


"\Delta T + \Delta U = W_f"


Let "h" be the maximum height the ball can reach. At this height ball is in rest and the change in kinetic energy is:


"\Delta T = 0-\dfrac{mv_0^2}{2 } = -\dfrac{mv_0^2}{2 }"

where "v_0" is a initial velocity.

The change in potential energy is:


"\Delta U = mgh - 0 = mgh"

The work done by friction force is:


"W_f = -F_f\cdot \dfrac{h}{\cos \theta}"

where "\theta = 30\degree." and "-" sign is because the force is directed in an opposite direction with respect to displacement.

The friction force for the rolling ball is:


"F_f = \dfrac{kmg}{R}"

where "k = 0.2 m" is a kinetic friction coefficient.

Combining it all together, obtain:


"\Delta T + \Delta U = W_f\\
-\dfrac{mv_0^2}{2 } + mgh = -\dfrac{kmg}{R}\cdot \dfrac{h}{\cos \theta}"

Expressing "h", get:


"h = \dfrac{v^2_0}{2g\left ( 1 + \frac{k}{R\cos\theta}\right )}"

Answer. "h = \dfrac{v^2_0}{2g\left ( 1 + \frac{k}{R\cos\theta}\right )}".